OpenStudy (anonymous):

How do you solve 30(1.2x) = 45 + 3x??

OpenStudy (anonymous):

want the answer or steps?

OpenStudy (anonymous):

both would be great

OpenStudy (anonymous):

30* 1.2x=45+3x multiply both sides by 10 10*30* 1.2x=10*45+10*3x refine 360x=30x+450 subtract 30x from both sides 360x-30x=30x+450-30x 330x+450 divide both sides by 330 330x/330=450/330 x=15/11 answer x=1.36364

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

no problem!